### v2

# Local Mixtures of Experts: Essentially Free Test-Time Training via Model Merging

Abstract
Mixture of expert (MoE) models are a promising approach to increasing model capacity without increasing inference cost, and are core components of many state-of-the-art language models. However, current MoE models typically use only few experts due to prohibitive training and inference cost. We propose Test-Time Model Merging (TTMM) which scales the MoE paradigm to an order of magnitude more experts and uses model merging to avoid almost any test-time overhead. We show that TTMM is an approximation of test-time training (TTT), which fine-tunes an expert model for each prediction task, i.e., prompt. TTT has recently been shown to significantly improve language models, but is computationally expensive. We find that performance of TTMM improves with more experts and approaches the performance of TTT. Moreover, we find that with a 1B parameter base model, TTMM is more than 100×100$\times$100 × faster than TTT at test-time by amortizing the cost of TTT at train-time. Thus, TTMM offers a promising cost-effective approach to scale test-time training.
Figure 1: Accuracy gains of TTMM when scaling test-time compute (left) and pre-training compute (right). The normalized perplexity between base model and TTT model is averaged across evaluation datasets and models. Left: We measure language modeling ability against the time to generate 100 tokens. TTMM approaches the performance of TTT without almost any test-time overhead when increasing the number of active experts. Ensembling rather than merging experts improves performance, but at a much higher test-time cost. Standard fine-tuning on the training data improves performance without any test-time overhead, but does not yield close to the performance of TTT. Right: We measure the performance of TTMM against the total number of experts. Whereas common MoE models use few experts, TTMM scales to an order of magnitude more experts, improving performance with more experts. Scaling the number of experts in TTMM does not affect training FLOPs or required GPU memory. [^2] Scaling the number of experts increases CPU host memory, which is cheaper than GPU memory. Training the individual experts is embarrassingly parallelizable.
1 Introduction
The standard paradigm of machine learning separates training and testing. In training, we learn a model by inductively extracting general rules from data, while in testing, we apply this model to new unseen data. We study an alternative transductive paradigm called test-time training (TTT, Sun et al., 2020) where we use a specific expert model for each prediction task. Variations of this paradigm have been studied since the inception of the machine learning field (Cleveland, 1979; Cleveland & Devlin, 1988; Atkeson et al., 1997; Bottou & Vapnik, 1992). More recently, fine-tuning large pre-trained neural networks at test-time has regained interest in computer vision (e.g., Jain & Learned-Miller, 2011; Shocher et al., 2018; Sun et al., 2020; Gandelsman et al., 2021; Ruiz et al., 2023) and language modeling (Krause et al., 2018; 2019; Hardt & Sun, 2024; Sun et al., 2024; Hübotter et al., 2025). Hardt & Sun (2024) and Hübotter et al. (2025) fine-tune a pre-trained language model on data related to the prompt, and show that this can significantly improve language modeling performance. Beyond pure language modeling, TTT has been instrumental in state-of-the-art approaches to abstract reasoning (Akyürek et al., 2025), video generation (Dalal et al., 2025), and also been shown to improve reasoning language models (Zuo et al., 2025; Simonds & Yoshiyama, 2025).
While TTT can significantly improve model performance, it comes with a high computational cost at test-time: TTT requires fine-tuning the model for every task, i.e., each prompt. We propose Test-Time Model Merging (TTMM) which amortizes the cost of TTT at train time:
  • At train-time, TTMM clusters the training data into many local neighborhoods and trains separate small expert LoRA adapters (Hu et al., 2022) for each cluster.
  • At test-time, TTMM dynamically selects a subset of LoRA adapters and merges their parameters to form a single task-specific model.
In our experiments, TTMM approaches the language modeling performance of TTT without incurring almost any significant compute or memory cost (cf. Footnote 2). Figure 2 illustrates the difference between TTT and TTMM.
Figure 2: Standard inductive language modeling pre-trains a single model on the dataset (shown in gray) which is then used for prediction. Test-Time Training (TTT) improves performance by fine-tuning the model at test-time on data related to the prompt, but is computationally expensive. Test-Time Model Merging (TTMM) fine-tunes the pre-trained model to many local neighborhoods at train-time, and merges the local models related to the prompt at test-time. Both TTT and TTMM target models specifically to the prompt (i.e., are transductive), yet, TTMM does not require expensive fine-tuning at test-time.
Another extensive body of work studies merging multiple models to combine distinct capabilities and knowledge (Yang et al., 2024b). We discuss the connection of TTMM to this literature on model merging and mixture of experts more extensively in Section 2. Model merging is typically used in a multitask setting with few “meta-tasks” (e.g., coding, math, law, etc.) and corresponding expert models. This multitask setting differs from the TTT setting where each individual prompt is considered its own task. Hence, TTMM trains many local models, even within a single meta-task (say for all kinds of different coding tasks). Then, for each prompt at test-time, TTMM merges the most related local models. As such, TTMM can be seen as bridging multitask model merging and TTT by dynamically constructing an expert model at test-time from an order of magnitude more meta-tasks than in multitask learning.
We contribute the following key findings, summarized in Footnote 2:
  1. 1.
    TTMM outperforms multitask model merging: We find that training and merging many expert models, even within a single domain such as “coding in Python”, outperforms training and merging few models specialized on broader tasks.
  2. 2.
    TTMM approaches performance of TTT without almost any test-time overhead: We discuss how to merge the parameters of the expert models at almost no additional compute or memory cost. We then evaluate TTMM on the Wikipedia and GitHub Python corpora and find that it achieves performance close to TTT, while matching the performance of ensembling without incurring its test-time overhead.
2 Related Work
We propose TTMM as an efficient amortization of TTT, which realizes comparable accuracy. We next discuss how TTMM can be seen as an extension to the literature on model selection and model merging. In Appendix A, we discuss the link to retrieval-augmented generation.
Multitask Model Merging.
Combining the predictions of multiple models has been a long-standing practice in machine learning, e.g., in the form of majority voting (Wang et al., 2023) or by ensembling of predictions (Dietterich, 2000; Wortsman et al., 2022), which has been used to avoid catastrophic forgetting (e.g., Bagatella et al., 2025). However, merging predictions of many neural networks is computationally inefficient since it requires a separate forward pass for each model. Instead, TTMM merges in parameter-space, and only requires a single forward pass of the merged model, resulting in a negligible increase of compute and memory cost compared to inferencing a single model. The idea of merging multiple models by averaging their parameters (not their predictions!) is grounded in the surprising phenomenon of mode-connectivity (Garipov et al., 2018; Wilson, 2025), which observes that modes in ensembles of models are connected by paths of small loss. This has motivated a variety of model merging techniques, including stochastic weight averaging (Izmailov et al., 2018) and model soups (Wortsman et al., 2022). Recently, model merging has been studied predominantly in the context of merging few expert models with the aim of retaining performance on their respective meta-tasks (Ilharco et al., 2023; Matena & Raffel, 2022; Jin et al., 2023; Ainsworth et al., 2023). Merging model parameters may lead to interference, which has been a key focus of recent work (Yadav et al., 2023; Yang et al., 2024c; Tang et al., 2024b; Daheim et al., 2024; Yu et al., 2024; Jung et al., 2024).
Dynamic Merging of Multitask Models.
Recent work in multitask model merging studies selecting merging coefficients dynamically depending on the prompt. Oh et al. (2025) select coefficients based on the entropy of the individual model’s predictions, which is computationally expensive since it requires forward passes of all networks. Instead, TTMM computes merging coefficients based on a measure of similarity between prompt and experts which is tuned on holdout data at train-time (Lu et al., 2024; Tang et al., 2024a; Cheng et al., 2024; Lai et al., 2025), and which incurs only negligible test-time overhead. In the tangentially related literature on black-box model fusion, dynamic model selection has been widely studied (e.g., Liu et al., 2021; Jiang et al., 2023; Mavromatis et al., 2024). However, in the setting of TTMM, we control data and architecture used for training the local models, which enables us to compute merging coefficients more efficiently than in black-box model fusion.
Clustering and Dynamic Model Selection.
While the above literature focuses on multitask learning with typically only a handful of expert models, TTMM clusters the training data into many local neighborhoods and trains a separate local model for each cluster. Like TTMM, Branch-Train-Merge (BTM, Li et al., 2022; Gururangan et al., 2023) clusters the training data into local neighborhoods, trains a separate local model for each cluster, and dynamically selects a subset of these models at test-time. However, unlike BTM, TTMM uses an order of magnitude more local models and merges their parameters instead of their predictions, leading to substantially faster inference. In concurrent work, Qiu et al. (2025) study TTMM in a continual learning setting.
Mixture of Experts.
The mixture of experts (MoE, Shazeer et al., 2017) paradigm increases model capacity without increasing inference cost by introducing conditional routing of inputs to subsets of learnable parameters (i.e., “experts”). Training MoE models is challenging due to the need for joint training of the experts and the router on substantial multitask data. Moreover, MoE models evaluate all active experts individually, which can still require substantial memory. This in turn implies that the number of experts is typically limited to a small number. Recent studies have focused on challenges such as reducing synchronization (Sukhbaatar et al., 2024; Zhang et al., 2024b; 2025), load balancing of experts (Clark et al., 2022; Zhou et al., 2022), and improving expert specialization (Dai et al., 2024). In contrast, TTMM approximates the MoE paradigm by embarrassingly parallelizing the training of the local models and merging model parameters at test-time, which permits scaling to many orders of magnitude more experts and which enables approximation of the “maximally local” TTT.
3 TTMM: Test-Time Model Merging
Our proposed method, TTMM, has two main stages: train-time and test-time. At train-time, TTMM clusters the training data into many local neighborhoods and trains a separate expert model for each cluster. Then at test-time, TTMM dynamically composes these expert models to form a single task-specific model. We discuss both stages in more detail below.
Figure 3: Illustration of TTMM: At train-time (Algorithm 1), TTMM clusters the training data (shown in gray) into many local neighborhoods and trains a separate expert model for each cluster. At test-time (Algorithm 2), TTMM dynamically selects a subset of expert models related to the prompt and merges their parameters to form a single task-specific model. The test-time stage is compute and memory efficient, using the previously trained experts.
We consider a domain $\mathcal{X}$ of token sequences and a training set $\mathcal{D}\subseteq\mathcal{X}$. We further assume that we have access to a pre-trained autoregressive language model $f(\boldsymbol{x};\boldsymbol{\theta})$ that maps token sequences $\boldsymbol{x}\in\mathcal{X}$ to probability distributions over the next token. Finally, we assume access to a sequence embedding model $\boldsymbol{\phi}(\boldsymbol{x})$ returning normalized embeddings. In our experiments, we use a mean-pooled sequence embedding model, but one could also use the last-layer token embeddings from the pre-trained language model (see Lu et al., 2024).
Algorithm 1 TTMM at train-time
1: Language model $f(\boldsymbol{x};\boldsymbol{\theta})$ with pre-trained weights $\boldsymbol{\theta}$; sequence embedding model $\boldsymbol{\phi}(\boldsymbol{x})$; training set $\mathcal{D}$; number of clusters $K$.
2:Cluster training set $\mathcal{D}$ into $K$ clusters $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$.
3:Train a LoRA adapter $\boldsymbol{\theta}_{k}$ on each cluster $\mathcal{D}_{k}$.
4:Compute cluster-specific embeddings:
$\boldsymbol{\phi}_{k}\leftarrow\mathrm{normalize}\left(\frac{1}{|\mathcal{D}_{k}|}\sum_{\boldsymbol{x}\in\mathcal{D}_{k}}\boldsymbol{\phi}(\boldsymbol{x})\right)$
Train-time: Clustering.
To cluster the training dataset, we use bisecting $k$-means, a hierarchical clustering algorithm which recursively applies $k$-means with $k=2$ to the cluster with the largest diameter (Jain, 2010). We use bisecting $k$-means since it is more efficient than $k$-means and tends to lead to more balanced clusters, however, any clustering algorithm can be used. We then train a small LoRA adapter on each cluster for one epoch. Finally, we average the embeddings of the sequences in each cluster to obtain cluster-specific embeddings (i.e., “centroids”) that represent the capabilities of the corresponding expert model. We renormalize the centroids to avoid biasing the embedding norm of larger clusters toward zero. Without the additional normalization, centroids of larger clusters have smaller norm despite $\boldsymbol{\phi}(\boldsymbol{x})$ being normalized.
Algorithm 2 TTMM at test-time
1: Prompt $\boldsymbol{x^{\star}}$ with embedding $\boldsymbol{\phi}^{\star}$. Language model $f(\boldsymbol{x};\boldsymbol{\theta})$ with pre-trained weights $\boldsymbol{\theta}$; expert models $\smash{\{(\boldsymbol{\theta}_{k},\boldsymbol{\phi}_{k})\}_{k=1}^{K}}$; temperature $\beta$ and sparsity $\tau$ (tuned on holdout data).
2:Compute cluster-specific merging coefficients:
$w_{k}\leftarrow\operatorname{sparse-softmax}_{\tau}\left(\frac{1}{\beta}\boldsymbol{\phi}_{k}^{\top}\boldsymbol{\phi}^{\star}\right)$
3:Merge into a single expert model:
$\boldsymbol{\theta}^{\star}\leftarrow\sum_{k=1}^{K}w_{k}\boldsymbol{\theta}_{k}$
4:Use $f(\boldsymbol{x^{\star}};\boldsymbol{\theta}^{\star})$ to generate the next token.
Test-time: Model Merging.
At test-time, given a prompt $\boldsymbol{x^{\star}}\in\mathcal{X}$ and its embedding $\boldsymbol{\phi}^{\star}\operatorname*{\,\dot{=}\,}\boldsymbol{\phi}(\boldsymbol{x^{\star}})$, we compute merging coefficients $w_{k}$ for each expert model $\boldsymbol{\theta}_{k}$ using a sparse cross-attention mechanism with sparsity parameter $\smash{\tau\in[0,\frac{1}{K})}$:
[…]
where $\operatorname{ReLU}(\boldsymbol{z})=\max\{0,\boldsymbol{z}\}$. This sparse softmax returns a sparse distribution over the experts, where experts with weight below $\tau$ are pruned, and the remaining weights are renormalized. Pruning experts with low weights improves test-time efficiency since less expert parameters need to be merged (see Figure 4 (right)), with minimal degradation of accuracy. The sparsity being smaller than $\nicefrac{{1}}{{K}}$ ensures that always at least one expert is selected. The model selection can easily be parallelized for all clusters and multiple queries by writing the computation as a cross-attention matrix product:
$\boldsymbol{\Theta}^{\star}\leftarrow\operatorname{sparse-softmax}_{\tau}\left(\frac{1}{\beta}\boldsymbol{\Phi}^{\star}\boldsymbol{\Phi}^{\top}\right)\boldsymbol{\Theta}$
with keys $\boldsymbol{\Phi}\operatorname*{\,\dot{=}\,}[\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, queries $\boldsymbol{\Phi}^{\star}\operatorname*{\,\dot{=}\,}[\boldsymbol{\phi}_{1}^{\star},\dots,\boldsymbol{\phi}_{n}^{\star}]^{\top}\in\mathbb{R}^{n\times d}$, and values $\boldsymbol{\Theta}\operatorname*{\,\dot{=}\,}[\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{K}]^{\top}\in\mathbb{R}^{K\times p}$, computing the collection of $n$ merged expert models $\boldsymbol{\Theta}^{\star}\in\mathbb{R}^{n\times p}$.
3.1 Latency
In contrast to MoE models, the number of experts in TTMM is typically too large for all experts to fit in GPU memory at once. For example, with Llama-3.2-1B as base model, each LoRA adapter has size approximately 172172172 MB, and with 100010001000 clusters, the expert models would require prohibitive 172172172 GB of GPU memory. Instead, we dynamically load active experts into GPU memory and merge their parameters before inference.
[…]
TTMM costs $\approx$20 tokens.
In Figure 4 (middle), we evaluate the practical test-time overhead of TTMM for language generation. TTMM incurs only a small fixed overhead which, when merging 101010 experts, is equal to generating around 202020 tokens. When generating more than 100100100 tokens, the overhead of TTMM compared to inference with the base model is negligible. In contrast, ensembling has a multiplicative overhead that grows with the number of active experts. Comparing to TTT, a single test-time gradient step with Llama-3.2-1B takes approximately 150150150 milliseconds (Hübotter et al., 2025). With 100100100 test-time gradient steps, the total test-time overhead (without any data selection and loading) is already around 151515 seconds. In contrast, TTMM with 101010 active experts has a constant overhead of around 115115115 milliseconds, which is a more than 125×125$\times$125 × speedup compared to TTT.
Balancing Latency and Performance.
The exact Pareto frontier of latency and performance varies depending on generation length. Footnote 2 plots the Pareto frontier for 100 generated tokens and Figure 4 (left) shows how the latency changes for other generation lengths. We do not show TTT in Figure 4 (left) since the constant overhead of TTT before generating any token is larger than the time to generate 1k tokens with TTMM (and 10 active experts).
3.2 Test-Time Model Merging approximates Test-Time Training
TTMM can be seen as an approximation of TTT that amortizes TTT’s compute cost at train-time. We formalize this connection by showing that in the limit of exponentially many experts, TTMM closely approximates TTT (Hardt & Sun, 2024), which fine-tunes the pre-trained model on the $N$ nearest neighbors to the prompt in the training data. To see the intuition behind this approximation, first consider the set of all subsets of $\mathcal{D}$ of size $N$. Suppose we pre-train an expert model on each of these subsets. Then, at test-time, we select the expert model whose training data embeddings have the shortest distance to the prompt embedding. Note that this is equivalent to TTT, without the need for fine-tuning at test-time. We have simply pre-computed all possible expert models at train-time, and at test-time fetch the right one. TTMM approximates this interpretation of TTT in three steps.
Approximation 1: Fewer amortized experts.
The number of pre-trained expert models in the above scenario is not practically feasible, as it grows exponentially with the training data. Thus, TTMM considers a more practical scenario with a smaller number of clusters where we do not pre-compute all possible TTT models. Our following proposition indicates that even an expert model different to the TTT model, yet trained on a cluster $\mathcal{D}^{\prime}$ sufficiently close to the prompt embedding, may still be a good approximation of TTT:
Informal Proposition 3.1 (Formalized in Proposition C.1).
Suppose the neural network $f(\boldsymbol{x};\boldsymbol{\theta})$ is $L$-Lipschitz in $\boldsymbol{\theta}$ and that the loss is $G$-smooth in $\boldsymbol{x}$, with any $L,G>0$. Further, suppose we adapt models via $T\geq 1$ steps of gradient descent with step size $\eta>0$ from a shared initialization $\boldsymbol{\theta}$. For any prompt $\boldsymbol{x^{\star}}\in\mathcal{X}$ and $N\geq 1$, let $\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}$ be the TTT model trained on the set $\mathcal{D}_{\boldsymbol{x^{\star}}}$ of the $N$ nearest neighbors to $\boldsymbol{\phi}(\boldsymbol{x^{\star}})$ in $\mathcal{D}$. Let $\boldsymbol{\theta}^{\prime}$ be the expert model trained on any $\mathcal{D}^{\prime}\subseteq\mathcal{D}$. Then, if $\mathcal{D}^{\prime}$ contains at least one nearest neighbor from $\mathcal{D}_{\boldsymbol{x^{\star}}}$,
$\left\|f(\boldsymbol{x^{\star}};\boldsymbol{\theta}_{\boldsymbol{x^{\star}}})-f(\boldsymbol{x^{\star}};\boldsymbol{\theta}^{\prime})\right\|\leq\eta TLG(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$ (1)
To understand the intuition behind this proposition, consider an expert model trained on a cluster $\mathcal{D}^{\prime}$ that contains parts of the local neighborhood of the prompt embedding (for this, the entire clustering must cover the dataset). Then, the above proposition shows that this expert model is a good approximation of the TTT model if the diameter of $\mathcal{D}^{\prime}$ is also small. Consequently, this motivates the approximation of TTT by TTMM with many expert models covering the entire dataset, each trained on a small cluster. Training expert models on this partitioning of the dataset does not cost more FLOPs or GPU memory than training a single model on the entire dataset. This contrasts with TTT, which considers a separate expert model for each possible recombination of the training data into subsets of size $N$.
Approximation 2: Summarizing experts with their centroids.
[…]
Approximation 3: Merging multiple experts.
As a final step, to compensate for the approximation error of the chosen expert model induced by steps (1) and (2), TTMM merges multiple expert models. Intuitively, and as visualized in Figure 3, this is motivated by multiple experts being able to cover distinct features of the prompt that may not all be captured by a single expert model. We find empirically that merging a small number of experts can improve model performance without incurring a significant test-time overhead (cf. Footnote 2). TTMM computes merging coefficients via cross-attention, which is equivalent to weighting the active experts with an RBF kernel: [^3] See Appendix B for a proof.
$\operatorname{softmax}\left(\frac{1}{\beta}\boldsymbol{\phi}_{k}^{\top}\boldsymbol{\phi}^{\star}\right)=\mathrm{normalize}\left(\exp\left(-\frac{1}{2\beta}\left\|\boldsymbol{\phi}^{\star}-\boldsymbol{\phi}_{k}\right\|_{2}^{2}\right)\right).$
Weighting data (not models!) by an RBF kernel has long been a common approach in machine learning and used, e.g., in kernel density estimation (Rosenblatt, 1956; Parzen, 1962), kernel regression (Nadaraya, 1964; Watson, 1964), RBF networks (Lowe & Broomhead, 1988; Moody & Darken, 1989), and local learning (e.g., Bottou & Vapnik, 1992; Atkeson et al., 1997). TTMM additionally sparsifies the selected models to minimize the test-time compute overhead due to loading and merging.
4 Results
Model Wikipedia GitHub (Python)
Merging Ensembling Merging Ensembling
Llama-3.2-1B 8.6748.6748.674 2.6112.6112.611
   + fine-tuning 7.8497.8497.849 2.5812.5812.581
   + TTMM (1 active expert) 7.6697.6697.669 2.5522.5522.552
   + TTMM (3 active experts) 7.571 7.579 2.519 2.512
   + TTMM (10 active experts) 7.510 7.509 2.492 2.464
   + TTT 7.559{7.559}7.559 $\underline{2.441}$
Qwen2.5-1.5B 8.570 2.335
   + fine-tuning 7.702 2.317
   + TTMM (1 active expert) 7.166 2.330
   + TTMM (3 active experts) 7.101 7.096 2.304 2.294
   + TTMM (10 active experts) 7.080 7.061 2.287 2.257
   + TTT 7.231 2.177
Table 1: Perplexity (lower is better) across evaluation datasets and base models with 100100100 experts. Bold numbers denote the best test-time efficient method, and underlined numbers denote the best overall. TTMM outperforms standard multitask fine-tuning of a single model.
We evaluate language modeling with causal language models. We report the perplexity of the models on holdout data from the Wikipedia (English, Wikimedia Foundation, 2025) and GitHub (Python code) corpora, comprising 6.4M and 7.2M documents, respectively. As base models, we use Llama-3.2-1B (Grattafiori et al., 2024) and Qwen2.5-1.5B (Yang et al., 2024a). For clustering and model selection, we use embeddings from the all-mpnet-base-v2 model (Reimers & Gurevych, 2019; Song et al., 2020). Following Hardt & Sun (2024) and Hübotter et al. (2025), we evaluate TTT by training a LoRA adapter for a single gradient step each on the 100100100 nearest neighbors to the prompt embedding, most to least similar. We also compare against a single model, which is fine-tuned for a single epoch on the respective corpus. This fine-tuned model can be considered an upper bound on the performance of approaches to multitask learning, since multitask learning typically trains one model per meta-task (i.e., per corpus) and merging models from multiple meta-tasks typically reduces performance on any individual meta-task.
Our code is available at https://github.com/rbertolissi/ttmerge. We share fine-tuned baselines and TTMM-MoEs on Huggingface.
Insight 1: TTMM approaches the accuracy of TTT at almost no test-time overhead.
We show in Footnote 2 and Table 1 that TTMM with 101010 active experts achieves close to the accuracy of TTT, and much higher accuracy than a single model that was fine-tuned on the respective corpus. Instead of merging experts in parameter space, we also evaluate merging experts in prediction space by ensembling the experts. We find that this yields marginally higher accuracy, yet it has a significantly larger test-time overhead. Overall, merging experts in parameter space achieves the best trade-off between accuracy and test-time efficiency. Interestingly, TTMM outperforms TTT on the Wikipedia dataset. This may be because TTMM can incorporate knowledge from a larger set of experts, each trained on a few thousands of data points. In contrast, TTT is only training on the 100 most related neighbors, which may miss some information if there are more than 100 related data points. In our view, the empirical difference between TTMM and TTT is an interesting direction for further study.
Figure 5: On each dataset, we evaluate the performance improvement of TTMM’s cross-attention mechanism compared to uniform weighting of active experts. Uniform weighting (dotted) selects the experts closest to the prompt embedding, and weights them equally. We see across merging (red) and ensembling (light blue) that the cross-attention mechanism outperforms uniform weighting significantly. Notably, on Wikipedia, combining multiple experts helps only when using cross-attention. The number of experts is $K=1000$.
Insight 2: Cross-attention is more effective than other merging coefficients.
In Figure 5, we evaluate perplexity with merging coefficients computed via cross-attention to uniform weighting of active experts. We find that cross-attention significantly outperforms uniform weighting. In particular, on the Wikipedia dataset, merging multiple experts does not help at all when weighting them uniformly. We also compare against merging coefficients, such as leveraging the logit-entropy of the experts (cf. Section D.3), but we do not find them to outperform TTMM’s cross-attention while being significantly more computationally expensive. We find that sparsity $\tau=0.01$ provides the best trade-off between accuracy and test-time efficiency, and we fix it in all other experiments. Furthermore, on our evaluation corpora, we need to select roughly 10 active experts to cover the local neighborhood of the prompt (cf. Figure 6 (right)), which is also when the performance of TTMM begins to saturate.
Refer to caption Refer to caption
[…]
Insight 3: Local experts outperform the base model in their local neighborhood.
We find in Figure 6 (left) that expert models perform significantly better on holdout data in their cluster than the base model or other expert models. This reaffirms the earlier findings of Hardt & Sun (2024) that TTT on local neighborhoods improves accuracy. We also find a large qualitative difference between the text generated by different experts and the base model, and we include examples in Appendix F.
Choosing the number of (active) experts.
We find, as summarized in Footnote 2, that TTMM with 100100100 experts significantly outperforms TTMM with 101010 experts or the multitask fine-tuned model (i.e., a single expert). This indicates that scaling the number of experts beyond the number of meta-tasks can increase accuracy substantially. Nevertheless, we also find that scaling the number of experts to 100010001000 does not further improve accuracy on our evaluation corpora. In the following, we discuss considerations for the number of experts in TTMM:
  • Number of experts ($K$): The total number of experts has almost no direct effect on latency. However, more experts require a large amount of static CPU memory. Further, increasing the number of experts also increases the separation of knowledge/skills between experts. This then requires more active experts to retain the same performance, which also increases latency and may lead to model interference due to merging. To prevent knowledge fragmentation, $K$ should be determined depending on the training set. In our experiments, choosing 100 experts is indicated by the “elbow method”, common in clustering (cf. Figure 6 (middle)).
  • Number of active experts ($N$): The number of active experts has the strongest effect on latency, since all active experts have to be loaded from CPU into GPU memory before inference. Our ablation in Figure 6 (right) indicates that with 101010 active experts, the “correct” expert is likely to be active, while latency remains minimal.
Evaluation on MMLU.
We additionally evaluate TTMM on MMLU (Hendrycks et al., 2021) as a non-perplexity task. We use (non-instruct) Llama-3.2-1B as base model, which we then fine-tune for one epoch on the MMLU training set. This fine-tuned model achieves approximately the same performance as instruction-tuned Llama-3.2-1B. We use the non-instruct model to avoid any confounding effect of RL-training or instruction-tuning on privileged data from the instruct model. As embedding model, we use Stella (stella_en_400M_v5) but obtain similar results with MPNet (cf. Section D.2). We train 100 experts on the MMLU training set, using the globally fine-tuned model as initialization. We summarize the results in Table 2 and find that TTMM slightly increases accuracy. A broader evaluation on other non-perplexity tasks and non-language domains is an important direction for future work.
Model Humanities Social Sciences STEM Other Overall
Fine-tuned 44.8 54.24 40.79 54.43 48.10
TTMM (1 expert) 45.25 54.92 40.98 54.30 48.41
TTMM (3 experts) 45.33 55.02 40.98 54.39 48.48
TTMM (10 experts) 45.63 55.15 41.39 54.55 48.74
TTMM (15 experts) 45.46 55.31 41.77 55.26 48.96
Table 2: Accuracy (in %) of TTMM on MMLU. Bold numbers denote the highest accuracy.
Ablations.
We perform a series of ablations, which we summarize in Appendix D. Notably, we find that (1) larger embedding models need not lead to improved performance of TTMM and that (2) TTMM performs similarly when fixing the number of experts as to when selecting the number of active experts dynamically via thresholding.
5 Discussion and Future Work
We propose Test-Time Model Merging (TTMM) which scales the mixture of experts paradigm to orders of magnitude more experts than in state-of-the-art multitask methods, leading to a substantial improvement in language modeling ability. We find that TTMM can achieve performance gain close to TTT by merging few active experts into a single model per query. TTMM achieves this with almost the same test-time cost as inference with the base model, and without increasing train-time FLOPs or GPU memory compared to multitask fine-tuning. A limitation of TTMM is the additional CPU memory required for storing all experts.
In achieving close to the performance of TTT without nearly any test-time overhead, TTMM opens up several exciting directions for future research. First, TTMM makes it feasible to re-select models online during generation or evaluation based on the growing prefix, which with TTT, is too expensive in most applications. Gururangan et al. (2023) have already shown that online re-selection can work with an ensemble selected from a smaller number of experts. Second, one may further improve the performance of TTMM using tools from the literature on model merging and MoEs aimed at avoiding interference between experts and improving the routing of data to experts, e.g., by fine-tuning the router on holdout data. TTMM can further be applied to settings with privileged information by adding or removing experts dynamically based on user access. Furthermore, it may be possible to leverage the large number of experts in TTMM for uncertainty estimation, e.g., by tracking also the variance of the merged parameters. Finally, many reasoning language models (called RLMs) such as DeepSeek R1 (Guo et al., 2025) are based on a MoE architecture, and TTT has been shown to be able to improve RLMs (Zuo et al., 2025). Since TTMM can be considered a scalable MoE architecture that approximates TTT, it would be particularly interesting to study whether TTMM can improve reasoning models.
Acknowledgments
This project was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943, and the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. Ido Hakimi was supported by an ETH AI Center Postdoctoral fellowship.
Appendices
[…]
0[2]
for name in lora_param_names: As = torch.stack([ d[name][”A”] for d in lora_dicts]) Bs = torch.stack([ d[name][”B”] for d in lora_dicts]) delta = torch.einsum( k i r, k r o -> i o”, Bs, As)
Figure 7: We merge LoRA adapters using PyTorch’s einsum, before applying the low-rank updates to the model parameters.
Appendix A Additional Related Work
Retrieval-Augmented Generation.
An alternative to specializing model parameters at test-time (i.e., TTT) is to instead specialize predictions by conditioning an autoregressive model directly on prompt-related data. This is called retrieval-augmented generation (RAG, Lewis et al., 2019; Khandelwal et al., 2020; Guu et al., 2020; Borgeaud et al., 2022) and motivated by the ability of LLMs to learn from context (Brown et al., 2020; Wei et al., 2022). Similarly to TTT, RAG leads to expensive inference since all local data needs to be processed at test-time. TTMM is more efficient by compressing local information into expert LoRA adapters at train-time. RAG is commonly used to endow LLMs with privileged information. Similarly, TTMM can include (& remove) information simply by adding (or removing) local experts.
Appendix B Equivalence of Cross-Attention and RBF Kernel
Assuming that all embeddings are normalized, we have
$\exp\left(-\frac{1}{2\beta}\left\|\boldsymbol{\phi}^{\star}-\boldsymbol{\phi}_{k}\right\|_{2}^{2}\right)$ $=\exp\left(-\frac{1}{\beta}+\frac{1}{\beta}\boldsymbol{\phi}_{k}^{\top}\boldsymbol{\phi}^{\star}\right)\propto\exp\left(\frac{1}{\beta}\boldsymbol{\phi}_{k}^{\top}\boldsymbol{\phi}^{\star}\right).$
Appendix C Proof of 3.1
Setup.
Let $f(\boldsymbol{x};\boldsymbol{\theta})$ be a neural network parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{p}$ and evaluated on input ${\boldsymbol{x}\in\mathcal{X}}$. We have a loss function $\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z})\in\mathbb{R}$ for data point $\boldsymbol{z}=(\boldsymbol{x},y)\in\mathcal{X}\times\mathcal{Y}$ [^4] This supervised data may be constructed via self-supervision, as in language modeling. Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{Y}$ be a dataset of input-output pairs. Given a prompt ${\boldsymbol{x^{\star}}\in\mathcal{X}}$ and any $N\geq 1$, we consider the set of $N$ nearest neighbors to $\boldsymbol{\phi}(\boldsymbol{x^{\star}})$ in $\mathcal{D}$ (with respect to the input embedding):
$\mathcal{D}_{\boldsymbol{x^{\star}}}=\{\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{N}\}\subseteq\mathcal{D}.$
Let $\mathcal{D}^{\prime}$ be any dataset of size $M\geq 1$:
$\mathcal{D}^{\prime}=\{\boldsymbol{z}_{1}^{\prime},\dots,\boldsymbol{z}_{M}^{\prime}\}\subseteq\mathcal{D}.$
We first analyze single-step gradient descent with step size $\eta>0$ from the same initialization $\boldsymbol{\theta}$:
$\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}=\boldsymbol{\theta}-\eta\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\,\boldsymbol{z}_{i})\right],\quad\boldsymbol{\theta}^{\prime}=\boldsymbol{\theta}-\eta\left[\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\,\boldsymbol{z}_{j}^{\prime})\right].$
Assumptions.
  1. (A1)
    Lipschitz-smoothness of the loss. There is a constant $G>0$ such that for all $\boldsymbol{z}=(\boldsymbol{x},y),\boldsymbol{z}^{\prime}=(\boldsymbol{x}^{\prime},y^{\prime})$ and all $\boldsymbol{\theta}$:
    $\bigl{\|}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\,\boldsymbol{z})\;-\;\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\,\boldsymbol{z}^{\prime})\bigr{\|}\;\leq\;G\,\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|.$
  2. (A2)
    Network Lipschitz in parameters. There is a constant $L>0$ such that for all $\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}$ and all $\boldsymbol{x}$:
    $\bigl{\|}f(\boldsymbol{x};\boldsymbol{\theta})\;-\;f(\boldsymbol{x};\boldsymbol{\theta}^{\prime})\bigr{\|}\;\leq\;L\,\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|.$
Proposition C.1.
Suppose (A1) and (A2) hold, and let $\boldsymbol{\theta}$ be any initialization. For any prompt ${\boldsymbol{x^{\star}}\in\mathcal{X}}$ and $N\geq 1$, let $\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}$ be the model trained via single-step gradient descent with step size ${\eta>0}$ on $\mathcal{D}_{\boldsymbol{x^{\star}}}$. Let $\boldsymbol{\theta}^{\prime}$ be the expert model trained on any $\mathcal{D}^{\prime}\subseteq\mathcal{D}$. Then, if $\mathcal{D}^{\prime}\cap\mathcal{D}_{\boldsymbol{x^{\star}}}\neq\emptyset$,
$\left\|f(\boldsymbol{x^{\star}};\boldsymbol{\theta}_{\boldsymbol{x^{\star}}})-f(\boldsymbol{x^{\star}};\boldsymbol{\theta}^{\prime})\right\|\leq\eta LG(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$ (2)
Proof.
We have
$\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}-\boldsymbol{\theta}^{\prime}$ $=-\eta\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{i})-\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{j}^{\prime})\right]$
$=-\eta\left[\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{i})-\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{j}^{\prime})\right]$
$=-\eta\left[\frac{1}{M}\sum_{j=1}^{M}\left(\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{i})-\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{j}^{\prime})\right)\right].$
Thus,
$\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}-\boldsymbol{\theta}^{\prime}\|$ $\leq\eta\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\|\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{i})-\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\mathcal{L}(\boldsymbol{\theta};\boldsymbol{z}_{j}^{\prime})\|$ (triangle inequality)
$\leq\eta G\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}^{\prime}\|.$ (Assumption A1)
Let $\boldsymbol{x}$ be any point in $\mathcal{D}^{\prime}\cap\mathcal{D}_{\boldsymbol{x^{\star}}}$. We have
$=\eta G\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\|\boldsymbol{x}_{i}-\boldsymbol{x}+\boldsymbol{x}-\boldsymbol{x}_{j}^{\prime}\|$
$\leq\eta G\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\left(\|\boldsymbol{x}_{i}-\boldsymbol{x}\|+\|\boldsymbol{x}-\boldsymbol{x}_{j}^{\prime}\|\right)$ (triangle inequality)
$\leq\eta G(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$
Finally,
$\|f(\boldsymbol{x^{\star}};\boldsymbol{\theta}_{\boldsymbol{x^{\star}}})-f(\boldsymbol{x^{\star}};\boldsymbol{\theta}^{\prime})\|$ $\leq L\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}-\boldsymbol{\theta}^{\prime}\|$ (Assumption A2)
$\leq\eta LG(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$
To complete the proof of 3.1, we simply observe that the bound $\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(1)}-\boldsymbol{\theta}^{\prime(1)}\|\leq\eta G(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime}))$ can be applied recursively from the shared initialization $\boldsymbol{\theta}^{(0)}$ for $T$ steps of gradient descent via the relation
$\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(t+1)}-\boldsymbol{\theta}^{\prime(t+1)}\|\leq\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(t)}-\boldsymbol{\theta}^{\prime(t)}\|+\eta G(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$
Unrolling this recursion proves 3.1 with an additional factor $T$ compared to Proposition C.1.
Appendix D Ablations
D.1 Fixing the Number of Active Experts
In Table 3, we ablate the choice of selecting the number of active experts dynamically per prompt depending on the threshold $\tau$ (cf. Algorithm 2) rather than specifying a fixed sparsity $N$. On our evaluated corpora, we find no significant difference in performance.
Model # Active Experts Wikipedia GitHub (Python)
TTMM Fixed TTMM Fixed
Llama-3.2-1B 1 7.669 7.669 2.552 2.552
3 7.571 7.580 2.519 2.505
10 7.510 7.509 2.492 2.491
Qwen2.5-1.5B 1 7.166 7.166 2.330 2.331
3 7.101 7.104 2.304 2.295
10 7.080 7.080 2.287 2.287
Table 3: Perplexity comparison (lower is better) between TTMM (dynamic expert routing) and fixed expert selection for Llama and Qwen across different numbers of active experts.
D.2 Different Embedding Models
In Table 4, we evaluate different embedding models on MMLU and find that TTMM outperforms the base model regardless of the choice of embedding model. Notably, the largest Qwen2-1.5B embedding model performs worse than the 100M MPNet model. A possible reason is that finding a decent clustering might be significantly easier than directly solving questions, since the former only requires a broad understanding of the meaning of some keywords while the latter requires some ability of reasoning. Stella (Zhang et al., 2024a) performs the best among the tested embedding models.
Model Humanities Social Sciences STEM Other Overall
Fine-tuned 44.80 54.24 40.79 54.43 48.10
all-mpnet-base-v2 45.95 54.76 40.56 54.72 48.61
stella_en_400M_v5 45.46 55.31 41.77 55.26 48.96
gte-Qwen2-1.5B-instruct 45.27 54.76 40.91 54.65 48.45
Table 4: MMLU accuracy (in %) comparing TTMM with different embedding models against the fine-tuned baseline. Bold indicates the highest accuracy.
D.3 Evaluation of Alternative Merging Coefficients
SIFT: Accounting for redundancy between experts.
We evaluate weighting active experts according to their estimated uncertainty reduction about the response to the prompt, as proposed by Hübotter et al. (2025). Building on transductive active learning (Hübotter et al., 2024), Hübotter et al. (2025) approximate the expected reduction in uncertainty about the prompt when conditioning on the $i$-th data point by $\smash{\sigma_{i-1}^{2}-\sigma_{i}^{2}}$ where $\smash{\sigma_{i}^{2}}$ estimates the uncertainty after seeing the $i$-th data point.
We use the same uncertainty estimates $\smash{\sigma_{i}^{2}}$ to weight the experts in TTMM, but instead of conditioning on individual data points represented by their embeddings, we condition on expert models represented by their centroids. For a given prompt, we first retrieve the $N$ nearest centroids and then reweight them according to:
$w_{k}\leftarrow\frac{\sigma_{k-1}^{2}-\sigma_{k}^{2}}{\sum_{k^{\prime}=1}^{N}\sigma_{k^{\prime}-1}^{2}-\sigma_{k^{\prime}}^{2}}=\frac{\sigma_{k-1}^{2}-\sigma_{k}^{2}}{\sigma_{0}^{2}-\sigma_{N}^{2}}.$
Given that embeddings are normalized, we have $\smash{\sigma_{0}^{2}=1}$.
Intuitively, these weights account for redundancy between clusters. Suppose we duplicate a cluster $m$ times, then the above weights downweigh each of the $m$ duplicates by a factor $\nicefrac{{1}}{{m}}$. The cross-attention mechanism in TTMM, on the other hand, would treat each duplicate cluster as if it were a distinct cluster, and hence increase the weight of the duplicated cluster by $m$.
Figure 8: We evaluate weighting experts according to SIFT’s uncertainty estimates, which take into account redundancy between clusters. Here, $\lambda$ is the hyperparameter of SIFT. As $\lambda\to\infty$, SIFT selects only the centroid closest to the prompt embedding, and as $\lambda\to 0$, SIFT weights all clusters uniformly.
We evaluate SIFT in Figure 8. We find that when we have control over the created experts (e.g., by clustering the data) as in TTMM, SIFT yields only a negligible improvmenet over the cross-attention mechanism. This likely stems from the fact that the centroids of the clusters are already well-separated, and hence the cross-attention mechanism is able to effectively weight the experts. We expect that in settings where we have less control over the experts, SIFT-weighting could provide a more significant improvement.
DaWin: Weighting experts by their logit-entropy.
Oh et al. (2025) propose DaWin, which weights experts by their logit-entropy. Let $p_{k}(\boldsymbol{x^{\star}})$ be the distribution over logits of expert $k$ for input $\boldsymbol{x^{\star}}$ with $H(p_{k}(\boldsymbol{x^{\star}}))$ denoting its entropy. Then, the weight of expert $k$ is given by
$w_{k}\leftarrow\operatorname{sparse-softmax}_{\tau}\left(-\frac{1}{\beta}H(p_{k}(\boldsymbol{x^{\star}}))\right).$
Note that these weights are equivalent to the weights obtained by the cross-attention mechanism in TTMM, up to the use of $\smash{-H(p_{k}(\boldsymbol{x^{\star}}))}$ to measure similarity as opposed to $\smash{\boldsymbol{\phi}_{k}^{\top}\boldsymbol{\phi}(\boldsymbol{x^{\star}})}$. As opposed to the inner product, which can be evaluated with very little overhead, computing the entropy requires a separate forward pass with each expert. This makes it as expensive as ensembling experts at test-time. In Figure 9, we compare the logit-entropy weighting to the inner product weighting of TTMM, and do not find a performance difference.
Figure 9: We evaluate ensembling models with inner product weighting (like in TTMM) and logit-entropy weighting (like in DaWin). We evaluate on Wikipedia with Llama-3.2-1B as base model. Solid lines are with sparsity $\tau=0.1$, translucent lines are with $\tau=0.01$.
Ablation against summarizing each expert by its clusters’ centroid.
As suggested in Section 3.2, we evaluate whether summarizing each expert by its clusters’ centroid meaningfully degenerates the performance of TTMM. We find that this appears not to be the case, as shown in Table 5.
Selection criterion Perplexity (lower is better)
Summarized by centroid $\left\|\boldsymbol{\phi}^{\star}-\boldsymbol{\phi}_{k}\right\|^{2}$ 7.669
Not summarized by centroid $\sum_{\boldsymbol{x}\in\mathcal{D}_{k}}\left\|\boldsymbol{\phi}^{\star}-\boldsymbol{\phi}(\boldsymbol{x})\right\|^{2}$ 7.633
Table 5: We evaluate the selection of experts, without approximating each expert by its clusters’ centroid as suggested in Section 3.2. The evaluation is on the Wikipedia corpus with Llama-3.2-1B as base model, $K=1000$ total experts and a single active expert.
Appendix E Experiment Details
Parameter Value
Number of clusters $K$ 100
Learning rate 2e-4
LoRA rank 64
LoRA alpha 16
Adam epsilon 1e-8
Adam beta (0.9, 0.999)
Weight decay 0.01
Batch size 4
Sparsity $\tau$ 0.01
Table 6: Hyperparameters.
Dataset Source
Wikipedia https://huggingface.co/datasets/wikimedia/wikipedia
GitHub (filtered to Python) https://huggingface.co/datasets/codeparrot/github-code-clean
Table 7: Datasets.
All models were trained on the first 1024 tokens of each training example. Testing was conducted on the full input length, with the context size limited to $\smash{2^{14}}$ tokens. For each dataset, the test sets were constructed by selecting a single random example from each cluster’s holdout set, ensuring that the evaluation approximately covers the data of all clusters.
The hyperparameters for TTT match those used to train the expert models, except for a learning rate of 5e-4. The models were trained for five epochs and the best checkpoint was selected for evaluation.
Unless noted otherwise, all ablations are with Llama-3.2-1B as the base model. Further, unless noted otherwise, we use “number of active experts” as an upper bound to the mean number of active experts across test examples.
Dataset size for expert training.
Both the Wikipedia and the GitHub (Python) corpora contain roughly 7M documents. Therefore, since each train document is at most 1024 tokens long, each expert is trained on at most $\approx$7M tokens if we train 100010001000 experts. This is a relatively small dataset size for training experts, compared to datasets used for training common MoE or multitask models.
Preventing Information Leakage.
The TTT setting (and also TTMM) requires specification of a prompt to obtain a task-specific model. In language modeling datasets that do not have the natural structure of prompt and response, it would be unfair to use the response as prompt. Using the response as prompt would invalidate the evaluation, as the test data is leaked to the model before its evaluation. To prevent this, we split individual data points into a prefix (prompt) and suffix (response) similar to Hardt & Sun (2024), use the prefix for obtaining the task-specific model, and evaluate the model on the suffix. Table 8 summarizes the choice of prefix per dataset. When not dynamically re-selecting models before the generation of each token (or after every $N$ tokens), the prefix must be sufficiently informative for the suffix. On Wikipedia, we find this to be around 50 tokens (cf., Figure 10).
Prefix length (in tokens)
Wikipedia 50
GitHub (Python) 200
Table 8: Overview of prefix lengths used per evaluation dataset.
Figure 10: On Wikipedia and with Llama-3.2-1B as base model, we evaluate different configurations of query prefix length (the number of tokens used as query) and evaluation prefix length (the number of tokens ignored for evaluation). Around 40 to 50 tokens, the performance gain of including more tokens in the query begins to diminish.
Appendix F Qualitative Examples
  1. 1.
    Clustering: To investigate the clustering, we use an LLM to summarize each cluster’s content with a title based on 10 data points of each cluster. Tables 9 and 10 list some of those cluster titles.
  2. 2.
    Expert Selection: Table 11 shows examples of selected experts for different queries.
  3. 3.
    Model Generation: We find that expert models generate distinctly different text. Table 12 gives examples of generations from empty strings, while Table 13 shows examples of generations from prompts. Within their subject of expertise, expert models tend to hallucinate less or at a later stage than the base model.
Cluster number Cluster title
118 Aviation History and Aviation-Related Topics
119 Historical Figures and Notable Americans
120 Historical Figures and Events in the United States
612 Honey Bees and their Scientific Classification
613 Small Mammals and Insects
614 Insect Taxonomy and Systematics
968 Swiss Geography and Mountains
969 Rivers of Germany
970 Historical Places and Landmarks in Europe
971 German Municipalities and Towns
Table 9: Wikipedia cluster titles
Cluster number Cluster title
17 Kivy GUI Framework
18 Tkinter GUI Framework Development
19 Multi-Touch Plugin for Alignak
427 Hidden Markov Model Implementation
428 Sparse Matrix Decomposition and Optimization
429 Logistic Regression with L2 Regularization
513 3D Printing and CAD Design
514 3D Mesh Generation and Manipulation Tools
979 Database Schema Definition and Migration Tools
980 MongoDB Database Schema Management
Table 10: GitHub (Python) cluster titles
Query input Closest expert 2nd-closest expert 3rd-closest expert
Basketball is a great sport (309) Women’s Basketball Competitions and Leagues (281) NBA and WNBA Teams and Seasons (311) Sports and Athletics
Albert Einstein was an astounding physicist (169) Historical Physicists and Scientists with Nobel prizes (104) Tufts University Alumni and Faculty in Science (88) British Scientists and Mathematicians
UNESCO World Heritage Site (716) Cities and Urban Landmarks (752) Asian Landmarks and Cultural Sites (906) Places in South Africa
Table 11: Example Wikipedia expert selection
Model Generated text
Base Model Ex 1. Question: What is the greatest common factor of 30 and 15? Answer: 15 Ex 2. Question: Design a Python function snippet to Check High Handling personal items: Sanitizing Frequently Touched Items for Analysis for Beginners. Implement if/else or switch/case statements to handle different conditions related to the Reliability. Provide detailed comments explaining your control flow and the reasoning behind each decision.
Expert (1): Politicians and Political Parties from Southeast Asia Ex 1. The 2018 South Korean by-elections was held in South Korea on 9 January 2018. The by-elections were held to fill two vacant seats in the National Assembly. The first by-election was held for the vacant seat in the constituency of Busan 1st District, and the second for the vacant seat in Seoul 6th District. Ex 2. The National Congress of Australia’s First Peoples (NCAFP) is an organisation that advocates for the interests of Aboriginal and Torres Strait Islander Australians in the Australian Parliament. It is a member of the Council of Australian Traditional Owner’s Organisations (CATOO).
Expert (280): American Football Teams and Players Ex 1. The 2017 season was the New England Patriots’ 48th in the National Football League (NFL), their 58th overall, their 51st playing home games at Gillette Stadium and their sixth under head coach Bill Belichick. The Patriots entered the season as the defending Super Bowl champions and were looking to defend their AFC East division title and make it to their third consecutive Super Bowl. Ex 2. The 2006 Houston Texans season was the franchise’s 7th season in the National Football League (NFL) and the 1st under head coach Gary Kubiak. They improved on their 4–12 record from 2005 and made the playoffs for the first time since 2003. They won the AFC South for the first time since 2002 and defeated the Indianapolis Colts in the Wild Card round 31–17.
Table 12: Generation from empty string inputs
Model Generated text
Base Model Ex 1. The politics in the US, the UK and elsewhere have been dominated by the “Brexit” vote to leave the European Union. While many people in the UK and elsewhere have been unhappy with the outcome, it is worth remembering that the vote was very close. The referendum was held on June 23rd, 2016, and the result was 52% in favour of leaving and 48% in favour of staying. Ex 2. Super Bowl 50: 5 Things to Know The 2016 Super Bowl will be played on February 7, 2016. It will be the 50th Super Bowl and the second to be played in the new 20,000 seat Levi’s Stadium in Santa Clara, California. Here are five things to know about the Super Bowl 50: 1. The Super Bowl is played on the first Sunday in February each year. The last Super Bowl was played on February 7, 2015.
Expert (1): Politicians and Political Parties from Southeast Asia Ex 1. The politics in the Cook Islands are characterised by a multi-party system and an electoral calendar dominated by the 2018 general election. The Cook Islands has had a democratic political system since 1965, with the Cook Islands being the first country in the Pacific to hold elections for the Parliament. Ex 2. Super Bowl XLIV was an American football game that was held on Sunday, February 7, 2010, at Superdome, New Orleans, Louisiana. The 2010 Super Bowl was the 45th edition of the American Football League Championship Game and the 46th overall Super Bowl.
Expert (280): American Football Teams and Players Ex 1. The politics in the 1990 NFL season involved the 1990 NFL Draft, the 1990 Pro Bowl, and the 1990 NFL season ending in controversy over the 1990 NFC Championship Game. The 1990 season was the first NFL season in which the league was divided into two conferences, with the NFC and AFC, and the first to end in controversy. Ex 2. Super Bowl LIV was an American football game played to determine the champion of the National Football League (NFL) for the 2019 season. The American Football Conference (AFC) champion Kansas City Chiefs defeated the National Football Conference (NFC) champion San Francisco 49ers, 31–20, on February 2, 2020, at Hard Rock Stadium in Miami Gardens, Florida.
Table 13: Generation from input strings, input is in bold